# Learning Pivoting Manipulation with Force and Vision Feedback Using Optimization-based Demonstrations

### v1

@@ Abstract
Non-prehensile manipulation is challenging due to complex contact interactions between objects, the environment, and robots. Model-based approaches can efficiently generate complex trajectories of robots and objects under contact constraints. However, they tend to be sensitive to model inaccuracies and require access to privileged information (e.g., object mass, size, pose), making them less suitable for novel objects. In contrast, learning-based approaches are typically more robust to modeling errors but require large amounts of data. In this paper, we bridge these two approaches to propose a framework for learning closed-loop pivoting manipulation. By leveraging computationally efficient Contact-Implicit Trajectory Optimization (CITO), we design demonstration-guided deep Reinforcement Learning (RL), leading to sample-efficient learning. We also present a sim-to-real transfer approach using a privileged training strategy, enabling the robot to perform pivoting manipulation using only proprioception, vision, and force sensing without access to privileged information. Our method is evaluated on several pivoting tasks, demonstrating that it can successfully perform sim-to-real transfer.
@@ 3 Method
[…]
@@
The non-privileged information $\mathbf{o}_{t}$ consists of the robot positions ${\mathbf{q}}_{t}^{r}\in\mathbb{R}^{2\times{N_{r}}}$, the binary robot contact signal ${d}_{t}\in\mathbb{Z}^{1}$, and the 2D contact forces ${\boldsymbol{\lambda}}_{t}^{r}\in\mathbb{R}^{2\times N_{r}}$ measured by force sensors mounted on the robots’ wrists. All observations are approximately normalized to lie in the range [1,1][-1,1][ - 1 , 1 ].
@@
$r_{\text{dyn}}=\alpha_{6}||{\mathbf{q}}_{t}^{r}-\phi(\mathbf{q}_{t}^{o})||^{2}+\alpha_{7}\arccos{\left(\frac{{\boldsymbol{\lambda}}_{t}^{r}\cdot\psi(\mathbf{q}_{t}^{o})}{||\boldsymbol{\lambda}_{t}^{r}||||\psi(\mathbf{q}_{t}^{o})||}\right)}+\alpha_{8}\mathbf{b}_{t}$ (4)
@@
Therefore, we define a student encoder that takes the history of the sensor observations, $[\mathbf{o}_{t-T},\cdots,\mathbf{o}_{t}]$, and the history of the segmentation features, $[\mathbf{s}_{t-T},\cdots,\mathbf{s}_{t}]$. Since $\mathbf{s}_{t}$ is high-dimensional, we first apply a Convolutional Neural Network (CNN) to compress the segmentation into a lower-dimensional feature representation $\mathbf{c}_{t}$. Using the temporal histories of $\mathbf{o}_{t}$ and $\mathbf{c}_{t}$, we use a Temporal Convolutional Network (TCN) [64] to estimate the privileged information. We train CNN and TCN jointly via supervised learning using datasets collected by rolling out the teacher policy in the simulator under domain randomization. The supervised learning objective is to minimize the following loss function:
$l=||\mathbf{p}_{t}-\tilde{\mathbf{p}}_{t}||^{2},$ (5)
@@ 4 Experiment Setup
[…]
Demonstration Setup. We use the method proposed in [57], randomizing object and environment parameters to generate diverse demonstrations. For all tasks, we randomize the mass of the object, the friction constant of the object and the environment, and the size of the object. For each task, we collect 500050005000 demonstrations, which can be computed within a few minutes using 303030 Intel i9-13900K CPU cores.
Teacher Policy Setup. We train the teacher policy in MuJoCo simulator [67] using robosuite [68] as a wrapper. The agent is trained using Soft Actor Critic (SAC) [69], implemented with tf2rl [70]. For SAC, we use Multi-Layer Perceptron (MLP) for both actor and critic networks. The simulation runs at 500 Hz, while the policy operates at 10 Hz. For each episode, we set the maximum episode length to 300300300 steps. Overall, training converges within 4 hours on a single NVIDIA RTX 4090. During training, we apply domain randomization over the objects’ mass and size, the friction constants of both the object and the environment, and the controller gains used in OSC within robosuite. Furthermore, we introduce sensor noise to both privileged information and sensor observations to account for the estimation errors from the student estimator during deployment.
Student Estimator Setup. We first rollout the trained teacher policy over 200020002000 episodes and collect a dataset containing ground-truth privileged information, sensor observations, and corresponding segmentation images (640×480 resolution) of the object using MuJoCo’s rendering functionality. During data collection, we augment the segmentation images by introducing noise, such as randomly flipping, translating, and rotating segmentation masks, to improve robustness. We then train the student estimator via behavior cloning, minimizing the loss function (5) over multiple epochs. We use $T=5$ step history of the observations for training corresponding to 0.5 second. Overall, training converges with 10 epochs (1 hour roughly), depending on the range of domain randomization.
@@ 5 Results
Figure 2: Learning curves for different RL training runs. Solid lines indicate average success rates, and shaded regions denote standard deviation across three different random seeds. Every 101010k step, the current policy is evaluated over 505050 episodes, and the success rate is plotted.
@@
[…]
@@
[…]
@@
Figure 4: Snapshots of successful pivoting manipulation in simulation and real-world.
@@ Appendix A CITO Details
[…]
@@
[…]
@@
[…]
@@ Appendix B Training Details in Simulation
$[\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{6},\alpha_{7},\alpha_{8}]$ [1, 0.075, 10, -1, -50, -50, -0.005, 5]
@@
The coordinate is illustrated in Fig. 5. In this work, we operate within the SE(2) group, restricting manipulation to the $y-z$ plane.
@@ B.1 Domain Randomization
[…]
@@
initial object position along $y$-axis +$\mathcal{U}(-0.015,0.015)$ m
initial robot position +$\mathcal{U}(-0.015,0.015)$ m
object position observation noise +$\mathcal{N}(0,0.015)$
robot position observation noise +$\mathcal{N}(0,0.00075)$
contact force observation noise +$\mathcal{N}(0,0.2)$
@@ B.2 Termination Conditions
[…]
@@
[…]
@@ C.1 Data Collection
Hole radius range [3,9][3,9][ 3 , 9 ] pixels
@@
Spatial Augmentation (Affine) Rotation range ±$\pm 2.5^{\circ}$
[…]
Scaling range [0.95, 1.05][0.95,\ 1.05][ 0.95 , 1.05 ]
@@ C.2 Student estimator training
Given the dataset collected in Section C.1, we train a student estimator composed of a CNN followed by a TCN. The CNN takes as input a binary segmentation mask of size 1×480×6401$\times$ 480$\times$ 6401 × 480 × 640 and consists of three convolutional layers with kernel sizes of (3,3,3)(3,3,3)( 3 , 3 , 3 ), and strides of (2,2,1)(2,2,1)( 2 , 2 , 1 ), and output channels of (16,32,64)(16,32,64)( 16 , 32 , 64 ), respectively. An adaptive average pooling layer reduces the spatial dimensions to $\times 8$ 8, followed by a fully connected layer that produces a 1×1281$\times$ 1281 × 128 feature vector. The TCN processes the temporal sequence of CNN features concatenated with proprioceptive and force features. It consists of three layers of 1D dilated causal convolutions, each with 128 channels and a kernel size of 2, and dilation rates of 1, 2, and 4. We consider two types of privileged information: time-invariant dynamics parameters (i.e., mass and size of the object), and time-varying values such as the object pose. To accommodate this distinction, the student estimator employs two separate fully connected layers—one for predicting the time-invariant variables and another for the time-varying privileged quantities (e.g., object pose). The output dimensions of each head match the corresponding target variables. We find that this separation leads to improved estimation performance.
@@ D.1 Effect of linear and quadratic reward terms during teacher policy training
Linear term No 88.1 ±0.21$\pm$ 0.21± 0.21
Quadratic term No 0.0 ±0.10$\pm$ 0.10± 0.10
Linear + Quadratic term No 88.9 ±0.20$\pm$ 0.20± 0.20
Linear term Yes 70.1 ±0.59$\pm$ 0.59± 0.59
Quadratic term Yes 0.0 ±0.71$\pm$ 0.71± 0.71
Linear + Quadratic term Yes 88.2 ±0.44$\pm$ 0.44± 0.44
@@ D.2 Pivoting with wall task without domain randomization
Figure 7: Learning curves for different RL training runs for pivoting-with-wall task. Solid lines indicate average success rates, and shaded regions denote standard deviation across three different random seeds. Every 101010k step, the current policy is evaluated over 505050 episodes, and the success rate is plotted.

### v2

@@ Abstract
Non-prehensile manipulation is challenging due to complex contact interactions between objects, the environment, and robots. Model-based approaches can efficiently generate complex trajectories of robots and objects under contact constraints. However, they tend to be sensitive to model inaccuracies and require access to privileged information (e.g., object mass, size, pose), making them less suitable for novel objects. In contrast, learning-based approaches are typically more robust to modeling errors but require large amounts of data. In this paper, we bridge these two approaches to propose a framework for learning closed-loop pivoting manipulation. By leveraging computationally efficient Contact-Implicit Trajectory Optimization (CITO), we design demonstration-guided deep Reinforcement Learning (RL), leading to sample-efficient learning. We also present a sim-to-real transfer approach using a privileged training strategy, enabling the robot to perform pivoting manipulation using only proprioception, vision, and force sensing without access to privileged information. Our method is evaluated on several pivoting tasks, demonstrating that it can successfully perform sim-to-real transfer. The overview of our method and the hardware experiments are shown here.
@@ 3 Method
s. t. $\text${s. t. }, $f_{\text{dyn}}\left(\bar{\mathbf{q}}_{t},\bar{\mathbf{q}}_{t+1},\dot{\bar{\mathbf{q}}}_{t},\bar{\mathbf{y}}_{t}\right)=\mathbf{0},g\left(\bar{\mathbf{q}}_{t},\dot{\bar{\mathbf{q}}}_{t},\bar{\mathbf{y}}_{t}\right)=\mathbf{0},$
@@
The non-privileged information $\mathbf{o}_{t}$ consists of the robot positions ${\mathbf{q}}_{t}^{r}\in\mathbb{R}^{2\times{N_{r}}}$, the binary robot contact signal ${d}_{t}\in\mathbb{Z}^{1}$, and the 2D contact forces ${\boldsymbol{\lambda}}_{t}^{r}\in\mathbb{R}^{2\times N_{r}}$ measured by force sensors mounted on the robots’ wrists. All observations are approximately normalized to lie in the range [1,1][-1,1].
@@
[…]
@@
Therefore, we define a student encoder that takes the history of the sensor observations, $\mathbf{o}_{t-T},\cdots,\mathbf{o}_{t}]$, and the history of the segmentation features, $\mathbf{s}_{t-T},\cdots,\mathbf{s}_{t}]$. Since $\mathbf{s}_{t}$ is high-dimensional, we first apply a Convolutional Neural Network (CNN) to compress the segmentation into a lower-dimensional feature representation $\mathbf{c}_{t}$. Using the temporal histories of $\mathbf{o}_{t}$ and $\mathbf{c}_{t}$, we use a Temporal Convolutional Network (TCN) [64] to estimate the privileged information. We train CNN and TCN jointly via supervised learning using datasets collected by rolling out the teacher policy in the simulator under domain randomization. The supervised learning objective is to minimize the following loss function:
[…]
@@ 4 Experiment Setup
We validate our framework across two distinct tasks (see Fig. 4): Pivoting with Wall: pivoting a box using an external wall, Pivoting without Wall: pivoting a box without relying on external support. For the latter task, the table surface must provide very high friction. In simulation, increasing the friction coefficient alone was insufficient to replicate the real-world behavior. As a workaround, we add a thin virtual wall of height $1\text{\,}\mathrm{mm}$ to simulate the effect of high-friction contact (see Fig. 4(b)). We define a trial as successful if the final orientation error satisfies $|\theta_{e}|\leq 0.087$ rad (i.e., $5\text{\,}\mathrm{\SIUnitSymbolDegree}$). We describe the setup for each module below, with additional details provided in the appendix.
Demonstration Setup. We use the method proposed in [57], randomizing object and environment parameters to generate diverse demonstrations. For all tasks, we randomize the mass of the object, the friction constant of the object and the environment, and the size of the object. For each task, we collect $5000$ demonstrations, which can be computed within a few minutes using $30$ Intel i9-13900K CPU cores.
Teacher Policy Setup. We train the teacher policy in MuJoCo simulator [67] using robosuite [68] as a wrapper. The agent is trained using Soft Actor Critic (SAC) [69], implemented with tf2rl [70]. For SAC, we use Multi-Layer Perceptron (MLP) for both actor and critic networks. The simulation runs at 500 Hz, while the policy operates at 10 Hz. For each episode, we set the maximum episode length to $300$ steps. Overall, training converges within 4 hours on a single NVIDIA RTX 4090. During training, we apply domain randomization over the objects’ mass and size, the friction constants of both the object and the environment, and the controller gains used in OSC within robosuite. Furthermore, we introduce sensor noise to both privileged information and sensor observations to account for the estimation errors from the student estimator during deployment.
Student Estimator Setup. We first rollout the trained teacher policy over $2000$ episodes and collect a dataset containing ground-truth privileged information, sensor observations, and corresponding segmentation images (640×480 resolution) of the object using MuJoCo’s rendering functionality. During data collection, we augment the segmentation images by introducing noise, such as randomly flipping, translating, and rotating segmentation masks, to improve robustness. We then train the student estimator via behavior cloning, minimizing the loss function (5) over multiple epochs. We use $T=5$ step history of the observations for training corresponding to 0.5 second. Overall, training converges with 10 epochs (1 hour roughly), depending on the range of domain randomization.
@@ 5 Results
Figure 2: Learning curves for different RL training runs. Solid lines indicate average success rates, and shaded regions denote standard deviation across three different random seeds. Every 1010k step, the current policy is evaluated over $50$ episodes, and the success rate is plotted.
@@
$50\text{\,}\mathrm{g}$ 2 / 3 3 / 3
$110\text{\,}\mathrm{g}$ 3 / 3 3 / 3
$300\text{\,}\mathrm{g}$ 0 / 3 3 / 3
How does the teacher policy’s performance vary with different demonstrations? For the pivoting-with-wall task, we deploy both kinematics- and dynamics-conditioned RL policies on a real system using a box of mass $110\text{\,}\mathrm{g}$. During deployment, we vary the mass values used as privileged information. Table 1 shows the success rates over three trials. We observe that dynamics-conditioned RL consistently outperforms kinematics-conditioned RL. While both policies are trained with access to privileged information, the dynamics-conditioned policy benefits from demonstrations that include contact force references. This enables the policy to learn physically grounded interaction behaviors during training, leading to greater robustness against variations in dynamic properties. In contrast, the kinematics-conditioned policy is trained with demonstrations that satisfy only geometric feasibility, making it more sensitive to changes in object properties. These results highlight the importance of dynamics-aware demonstrations in contact-rich manipulation tasks.
How robust is the learned policy compared to MPC? We compare the robustness of a dynamics-conditioned RL policy against an MPC controller on the real-world pivoting-with-wall task. The true object length is $0.16\text{\,}\mathrm{m}$, and we introduce intentional mismatches in the assumed object length during deployment. For example, a $5\text{\,}\mathrm{mm}$\$ offset means that the actual size of the box is shorter than what the controllers expect. As shown in Table 2, both MPC and RL succeed when the actual object is longer than expected ($5\text{\,}\mathrm{mm}$\$), as the contact with the wall is still maintained. However, when the actual object is shorter than expected ($5\text{\,}\mathrm{mm}$\$), MPC fails completely, while RL remains successful. This suggests that the learned policy exhibits greater tolerance to moderate discrepancies in privileged information. At larger mismatches ($10\text{\,}\mathrm{mm}$\$), even RL fails. These results highlight the importance of accurate privileged information during deployment and motivate us to develop reliable estimators.
@@
$5\text{\,}\mathrm{mm}$\$ 3 / 3 3 / 3
$5\text{\,}\mathrm{mm}$\$ 0 / 3 3 / 3
$10\text{\,}\mathrm{mm}$\$ 0 / 3 0 / 3
@@
Figure 4: Snapshots of successful pivoting manipulation in simulation and real-world. All hardware experiment videos are found at https://youtu.be/akjGDgfwLbM?si=0umnlPv-PMCTk_6F.
@@ Appendix A CITO Details
s. t. $\text${s. t. }, $h_{1}\left(\bar{\mathbf{q}}_{t}^{o},\bar{\mathbf{q}}_{t+1}^{o},\dot{\bar{\mathbf{q}}}_{t}^{o}\right)=\mathbf{0},$
@@
s. t. $\text${s. t. }, $h_{2}\left(\bar{\mathbf{q}}_{t}^{r},\bar{\mathbf{q}}_{t+1}^{r},\dot{\bar{\mathbf{q}}}_{t}^{r},\bar{\mathbf{y}}_{t}\right)=\mathbf{0},$
@@
s. t. $\text${s. t. }, $h_{3}\left(\bar{\mathbf{q}}_{t}^{r},\bar{\mathbf{q}}_{t+1}^{r},\dot{\bar{\mathbf{q}}}_{t}^{r},\bar{\mathbf{y}}_{t}\right)=\mathbf{0},$
@@ Appendix B Training Details in Simulation
$\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},\alpha_{5},\alpha_{6},\alpha_{7},\alpha_{8}] [$1, 0.075, 10, -1, -50, -50, -0.005, 5]
@@
The coordinate is illustrated in Fig. 5. In this work, we operate within the SE(2) group, restricting manipulation to the yzy-z plane.
@@ B.1 Domain Randomization
Table 5: Dynamics randomization and sensor noise. $\mathcal{N}(\mu,\sigma)$ denotes a Gaussian distribution with mean $\mu$ and standard deviation $\sigma$, and $\mathcal{U}(a,b)$ denotes a uniform distribution over the interval of $[a,b]$. A ++ symbol indicates that the sampled noise is added to the original parameter value.
@@
initial object position along $y$-axis $\mathcal{U}(-0.015,0.015)$ m
initial robot position $\mathcal{U}(-0.015,0.015)$ m
object position observation noise $\mathcal{N}(0,0.015)$
robot position observation noise $\mathcal{N}(0,0.00075)$
contact force observation noise $\mathcal{N}(0,0.2)$
@@ B.2 Termination Conditions
    Successful task completion: A trial is considered successful if the final orientation error satisfies $|\theta_{e}|\leq 0.087$ radians (i.e., $5\text{\,}\mathrm{\SIUnitSymbolDegree}$).
@@
    Significant deviation from the SE(2) plane: If the object’s $x$-position $p_{x}$ deviates by more than $0.05\text{\,}\mathrm{m}$ from its initial value, i.e., $|p_{x}-p_{x}(t=0)|\geq$0.05\text{\,}\mathrm{m}$\$, or if the $z$-position drops below the table surface, $p_{z}\leq p_{z}^{\text{table}}$, the episode is terminated and a penalty of -100 is applied.
@@ C.1 Data Collection
Hole radius range $[3,9]$ pixels
@@
Spatial Augmentation (Affine) Rotation range $\pm 2.5^{\circ}$
Translation range $\pm 7.5\%$
Scaling range [0.95, 1.05][0.95,\ 1.05]
@@ C.2 Student estimator training
Given the dataset collected in Section C.1, we train a student estimator composed of a CNN followed by a TCN. The CNN takes as input a binary segmentation mask of size $1\times 480\times 640$ and consists of three convolutional layers with kernel sizes of $(3,3,3)$, and strides of $(2,2,1)$, and output channels of $(16,32,64)$, respectively. An adaptive average pooling layer reduces the spatial dimensions to $8\times 8$, followed by a fully connected layer that produces a $1\times 128$ feature vector. The TCN processes the temporal sequence of CNN features concatenated with proprioceptive and force features. It consists of three layers of 1D dilated causal convolutions, each with 128 channels and a kernel size of 2, and dilation rates of 1, 2, and 4. We consider two types of privileged information: time-invariant dynamics parameters (i.e., mass and size of the object), and time-varying values such as the object pose. To accommodate this distinction, the student estimator employs two separate fully connected layers—one for predicting the time-invariant variables and another for the time-varying privileged quantities (e.g., object pose). The output dimensions of each head match the corresponding target variables. We find that this separation leads to improved estimation performance.
@@ D.1 Effect of linear and quadratic reward terms during teacher policy training
Linear term No 88.1 $\pm 0.21$
Quadratic term No 0.0 $\pm 0.10$
Linear + Quadratic term No 88.9 $\pm 0.20$
Linear term Yes 70.1 $\pm 0.59$
Quadratic term Yes 0.0 $\pm 0.71$
Linear + Quadratic term Yes 88.2 $\pm 0.44$
@@ D.2 Pivoting with wall task without domain randomization
Figure 7: Learning curves for different RL training runs for pivoting-with-wall task. Solid lines indicate average success rates, and shaded regions denote standard deviation across three different random seeds. Every 1010k step, the current policy is evaluated over $50$ episodes, and the success rate is plotted.